Assume 2 kg of hydrogen (Item B 28) are initially at 40 atm, 100 K. (a) If the hydrogen changes pressure to 1 atm without a change of temperature, find the changes of enthalpy and entropy. (b) If the hydrogen changes temperature to 35 K with no change of entropy, find the change of pressure and enthalpy.

1
Expert's answer
2021-11-16T13:44:00-0500

"m = 2\textsf{ kg}\\
P_1 = 40\textsf{ atm}\\
T_1 = 100\textsf{ K}\\
P_2= 1\textsf{ atm}\\
T_2 = 100\textsf{ K}\\
∆P = 39\textsf{ atm}"


Change of enthalpy "(ΔH) = ΔE + Δ(PV)= 0 + mRT\ln\dfrac{P_1}{P_2} = 2 × 8.314 × 100 × \ln\dfrac{40}1=6133.9\ J"


Change of entropy = "\dfrac{∆Q}T = 0"


Change of enthalpy = ΔH = ΔE + Δ(PV)"."